Work problems:
1. Two weather satellites working together, each starting at opposite directions from the North Pole, can take necessary photographs to map the entire earth sky in 20 hours. After they hadworked for 12 hours, the first satellite developed communications problem with its earth controlling base, and the second had to finish the job, requiring 1 day additional load time. How long would it take each satellite alone to do the weather-mapping task?
2.Joshua estimates that he could paint Mrs. De Leon's house in 8 days, David could also do the work in 9 days. How long would it take both of them to finish the job if they worked together?

Motion Problems:
1. A ferry boat cruises between two islands that are 40 kilometers apart. Travelling with a slow ocean current of 10kph when going and againts it when returning, the round trip takes 20 minutes more time running. What is the speed of the ferry boat in still ocoan water? How long did the trip take each way?

That's my assignment. Can you help me answer this? please. I need it tomorrow. Thank You.

Answer by mananth(16949)

Satellite I and satellite II take 20 hours to do the job
They work together for 12 hours
so they finsish 12/20 of the job
= 3/5
Balance job is 1- 3/5 = 2/5
2/5 of the job the satellite II does in 24 hours
so in 1 hour it does 2/24*5 of the job = 2/120 = 1/60
Together they can do the job in 20 hours
Satellite II does 1/60 of the job in 1 hour

so Satellite I does 1/20 -1/60 of the job in 1 hour
=2/60 = 1/30 of the job in 1hour.
Satellite takes 60 hours to do the job alone
Satellite I takes 30 hours to do the job alone.

Joshua does 1/8 of the job in 1 hour
David does 1/9 of the job in 1 hour
so together they do 1/8 +1/9 of the job in 1 hour
=17/72 of the job in 1 hour
so they take 72/17 hours to complete the job together.
4.23 hours

speed of boat be x
current = 10 kph
speed with current = x+10
speed against current = x-10
Return trip takes 20 minutes more
time going =time return + 1/3 hours
OR
Time return - time going = 1/3
Time = distance / speed
40/(x-10) -40/(x+10)= 1/3
LCD = 3(x+10)(x-10)
3*40(x+10)-3*40(x-10)= (x+10)(x-10)
120x+1200 -120x+1200=x^2-100
x^2= 2500
x= 25
speed of boat = 25 kph
m.ananth@hotmail.ca